Statement of Cash Flows-Indirect Method The comparative balance sheet of Olson-Jones Industries Inc. for December 31, 20Y2 anNet cash flow from operating activities Cash flows from (used for) investing activities: Cash from sale of land 23 Cash usedStatement of Cash Flows-Indirect Method The comparative balance sheet of Merrick Equipment Co. for December 31, 2049 and 2048Total liabilities and stockholders equity $930,190 $805,540 Additional data obtained from an examination of the accounts inChanges in current operating assets and liabilities: Net cash flow from operating activities Cash flows from (used for) inves
Answer #1
Cash Flow Statement
Indirect Method
Cash flow from Operating Activities
Net Income $        64,950.00
Adjustments
Depreciation $      7,110.00 =65110-58000
Loss on sale of Investments $      9,420.00 =94250-84830
Change In current assets & Liabilities
Increase in Accounts Receivable $ 7,130.00 =82170-89300
Increase in inventories $ 8,800.00 =243270-252070
Increase in Accounts payable $      9,670.00 =168360-158690
Decrease in Accrued Expenses Payable $ 4,200.00 =16740-20940
Total Adjustments $          6,070.00
Cash from operating activities $        71,020.00
Cash flow from Investing Activities
Sale of Investment $        84,830.00
Cash used for Purchase of Land $ 1,29,300.00 =0-129300
Cash used for Purchase of Equipment $ 63,050.00 =215080-278130
Net Cash Flow used for investing activities $ 1,07,520.00
Cash flow from Financing Activities
Issue of stock $        90,040.00 =50230-39470+188830-109550
Cash used for Dividends $ 35,810.00 =-(37860+7250-9300)
Cash from financing activities $        54,230.00
Net Increase in cash $        17,730.00
Opening Balance of Cash $     2,28,770.00
Closing Balance of Cash $     2,46,500.00